the perimeter of a rectangle is 200ft how do u find the lengh if the width is 20

Answer by mananth(16946)

Perimeter P = 2*(L+W)
200=2*(l+20)
divide by 2
200/2 = l+20
100=l+20
-20
l+20-20=100-20
l=80